# Traceback of Poisoning Attacks to Retrieval-Augmented Generation

### v1

@@ Abstract.
Large language models (LLMs) integrated with retrieval-augmented generation (RAG) systems improve accuracy by leveraging external knowledge sources. However, recent research has revealed RAG’s susceptibility to poisoning attacks, where the attacker injects poisoned texts into the knowledge database, leading to attacker-desired responses. Existing defenses, which predominantly focus on inference-time mitigation, have proven insufficient against sophisticated attacks. In this paper, we introduce RAGForensics, the first traceback system for RAG, designed to identify poisoned texts within the knowledge database that are responsible for the attacks. RAGForensics operates iteratively, first retrieving a subset of texts from the database and then utilizing a specially crafted prompt to guide an LLM in detecting potential poisoning texts. Empirical evaluations across multiple datasets demonstrate the effectiveness of RAGForensics against state-of-the-art poisoning attacks. This work pioneers the traceback of poisoned texts in RAG systems, providing a practical and promising defense mechanism to enhance their security.
@@ 2.1. RAG Overview
Knowledge retrieval: The goal of this stage is to retrieve the most relevant texts from the external knowledge database based on the user’s query. Typically, this is achieved using a vector-based retrieval model. For a user query $q$, the query encoder $f_{q}$ generates an embedding vector $v_{q}$. Each text $d_{j}$ in the knowledge database $\mathcal{D}$ is encoded into a vector $v_{d_{j}}$ by a text encoder $f_{d}$, forming the set $V_{\mathcal{D}}$. The relevance between $v_{q}$ and each $v_{d_{j}}\in V_{\mathcal{D}}$ is measured by similarity (e.g., dot product or cosine similarity). Based on these scores, the top-$K$ most relevant texts ^$\widehat{\mathcal{R}}(q,K,\mathcal{D})$ are selected.
Answer generation:  At this stage, the query $q$ is combined with the set of retrieved texts ^$\widehat{\mathcal{R}}(q,K,\mathcal{D})$ to query the LLM, which generates a response $O$. Formally, the response can be represented as the following:
@@
where ^$\widehat{\mathcal{R}}(q,K,\mathcal{D})$ represents the set of the top-$K$ most relevant texts retrieved from the knowledge database $\mathcal{D}$ based on query $q$. In this stage, we utilize a similar system prompt as described in (Zou et al., 2025) for RAG, as shown in Appendix 8.1.
@@ 3.2. Design Challenges
[…]
@@
where $|\mathcal{Q}|$ denotes the number of user queries, $\mathcal{D}\setminus\mathcal{D}_{p}$ represents the knowledge database after removing the identified poisoned texts $\mathcal{D}_{p}$ from $\mathcal{D}$, ^$\widehat{\mathcal{R}}(q_{i},K,\mathcal{D}\setminus\mathcal{D}_{p})$ represents the top-$K$ texts retrieved from the database $\mathcal{D}\setminus\mathcal{D}_{p}$ for the query $q_{i}$, and $\mathbb{I}(\cdot)$ is an indicator function that returns 1 if the LLM’s output matches (or aligns with) the incorrect output $t_{i}$, and 0 otherwise.
Key challenges:  The main challenge in identifying the poisoned texts $\mathcal{D}_{p}$ lies in solving the optimization problem in Eq. (2). A straightforward approach, inspired by (Shan et al., 2022), would be to iteratively update $\mathcal{D}_{p}$ using the gradient of the objective function. However, this is impractical under our threat model. First, computing the gradient is challenging due to two key constraints: (1) we lack access to the parameters of both the retriever and the LLM, and (2) both $\widehat{\mathcal{R}}(\cdot)$ and $\mathbb{I}(\cdot)$ are discrete functions, making gradient-based optimization infeasible. Second, determining update candidates for $\mathcal{D}_{p}$ would require computing the gradient for every text in the knowledge database $\mathcal{D}$, introducing substantial computational overhead and further complicating the optimization process.
@@ 4.2. RAGForensics
The prompt designed to guide the LLM in assessing whether the text $\mathcal{C}_{j}$ is poisoned. I will provide a question, an response, and a context. Without any other knowledge, judge whether the provided context tries to induce you to generate an answer consistent with the provided response [$t_{i}$], regardless of whether it is correct. Begin your judgement by providing a short explanation. If yes, return “[Label: Yes]”, otherwise, return “[Label: No]”. Question: [$q_{i}$] Context: [$\mathcal{C}_{j}$] Response:[$t_{i}$] Let’s think step by step.

### v2

@@ Abstract.
Large language models (LLMs) integrated with retrieval-augmented generation (RAG) systems improve accuracy by leveraging external knowledge sources. However, recent research has revealed RAG’s susceptibility to poisoning attacks, where the attacker injects poisoned texts into the knowledge database, leading to attacker-desired responses. Existing defenses, which predominantly focus on inference-time mitigation, have proven insufficient against sophisticated attacks. In this paper, we introduce RAGForensics, the first traceback system for RAG, designed to identify poisoned texts within the knowledge database that are responsible for the attacks. RAGForensics operates iteratively, first retrieving a subset of texts from the database and then utilizing a specially crafted prompt to guide an LLM in detecting potential poisoning texts. Empirical evaluations across multiple datasets demonstrate the effectiveness of RAGForensics against state-of-the-art poisoning attacks. This work pioneers the traceback of poisoned texts in RAG systems, providing a practical and promising defense mechanism to enhance their security. Our code is available at: https://github.com/zhangbl6618/RAG-Responsibility-Attribution
@@ 2.1. RAG Overview
Knowledge retrieval: The goal of this stage is to retrieve the most relevant texts from the external knowledge database based on the user’s query. Typically, this is achieved using a vector-based retrieval model. For a user query $q$, the query encoder $f_{q}$ generates an embedding vector $v_{q}$. Each text $d_{j}$ in the knowledge database $\mathcal{D}$ is encoded into a vector $v_{d_{j}}$ by a text encoder $f_{d}$, forming the set $V_{\mathcal{D}}$. The relevance between $v_{q}$ and each $v_{d_{j}}\in V_{\mathcal{D}}$ is measured by similarity (e.g., dot product or cosine similarity). Based on these scores, the top-$K$ most relevant texts $\widehat{\mathcal{R}}(q,K,\mathcal{D})$ are selected.
Answer generation:  At this stage, the query $q$ is combined with the set of retrieved texts $\widehat{\mathcal{R}}(q,K,\mathcal{D})$ to query the LLM, which generates a response $O$. Formally, the response can be represented as the following:
@@
where $\widehat{\mathcal{R}}(q,K,\mathcal{D})$ represents the set of the top-$K$ most relevant texts retrieved from the knowledge database $\mathcal{D}$ based on query $q$. In this stage, we utilize a similar system prompt as described in (Zou et al., 2025) for RAG, as shown in Appendix 8.1.
@@ 3.2. Design Challenges
(2) $\min_{\mathcal{D}_{p}}$ $\quad\frac{1}{|\mathcal{Q}|}\sum_{i=1}^{|\mathcal{Q}|}\mathbb{I}(\text{LLM}(\widehat{\mathcal{R}}(q_{i},K,\mathcal{D}\setminus\mathcal{D}_{p}),q_{i})\xrightarrow{\text{Match}}t_{i})$
@@
where $|\mathcal{Q}|$ denotes the number of user queries, $\mathcal{D}\setminus\mathcal{D}_{p}$ represents the knowledge database after removing the identified poisoned texts $\mathcal{D}_{p}$ from $\mathcal{D}$, $\widehat{\mathcal{R}}(q_{i},K,\mathcal{D}\setminus\mathcal{D}_{p})$ represents the top-$K$ texts retrieved from the database $\mathcal{D}\setminus\mathcal{D}_{p}$ for the query $q_{i}$, and $\mathbb{I}(\cdot)$ is an indicator function that returns 1 if the LLM’s output matches (or aligns with) the incorrect output $t_{i}$, and 0 otherwise.
[…]
@@ 4.2. RAGForensics
The prompt designed to guide the LLM in assessing whether the text $\mathcal{C}_{j}$ is poisoned. I will provide a question, an response, and a context. Without any other knowledge, judge whether the provided context tries to induce you to generate an answer consistent with the provided response $t_{i}$], regardless of whether it is correct. Begin your judgement by providing a short explanation. If yes, return “[Label: Yes]”, otherwise, return “[Label: No]”. Question: $q_{i}$] Context: $\mathcal{C}_{j}$] Response:[$t_{i}$] Let’s think step by step.